OpenStudy (anonymous):

find the slope of the line determined by the equation 2x=y

OpenStudy (anonymous):

general equation of a line is y = mx+b where m is the slope and b is the y-intercept. from your given, we only have y = 2x, or we should say y = 2x+0 to make it agree with the equation of a line. so, what could be your slope?

OpenStudy (anonymous):

tell weather the following equation determines y to be a function of x: y2= x + 3

Directrix (directrix):

y2= x + 3 ---------- x = y^2 - 3 ==> x is a function of y.